OpenStudy (anonymous):

solve the equation 2 cos^2 x = 3sinx, for 0^degree < x < 360^degree

OpenStudy (anonymous):

Use \[ \cos^2x=1-\sin^2x \]

OpenStudy (anonymous):

\[ 2 \cos^2 x = 3\sin x \]Becomes \[ 2 -2\sin^2 x = 3\sin x\] Now let \(u=\sin x\)\[ 2 -2u^2 = 3u \]This is just a quadratic equation. Solve for \(u\) and remember \(x=\sin^{-1}u\)

OpenStudy (anonymous):

thanks! really helps me alot :)

OpenStudy (anonymous):

Good luck!